# RadioDiff: An Effective Generative Diffusion Model for Sampling-Free Dynamic Radio Map Construction

### v1

@@ I Introduction
In wireless networks, pathloss quantifies the attenuation of signal strength between a pair of sender and receiver caused by free-space propagation loss and interactions of radio waves with obstacles [1, 2, 3], which is critical to be measured for wireless resource management [4, 5, 6]. Traditionally, the pathloss measurement usually depends on pilot transmission and signal processing [7]. However, the dramatic increase of network nodes and antennas has led to the challenge of estimating high-dimensional channels [8], resulting in significant costs in training and feedback overhead, as well as signal processing complexity. This issue further deteriorates in high-mobility scenarios with short channel coherence times for low-latency applications [9, 10]. Meanwhile, the upcoming 6G networks will introduce a large variety of node types, including passive equipment such as intelligent reflective surfaces (IRS) [11], which cannot actively transmit pilots or engage in digital signal processing to measure the pathloss.
@@
The fundamental reason for the above challenges is the misalignment between the generative problem attributes of the RM construction and the discriminant methodology. RM construction shows the characteristics of generative problems in both data features and NN training methods, while the existing sample-free NN-based methods use discriminant methods to construct RM, which inevitably limits the construction performance of RM. Specifically, in terms of data features, neither the value nor location of the elements of the pathloss to be predicted exist in the input environmental data, thus NN is required to generate pathloss from the raw data. Moreover, since the elements in the RM are not discrete values, it is almost impossible to construct the RM by dividing a finite number of hyperplanes to classify the elements in the environmental data as a discriminative problem [23]. Therefore, the use of discriminant training methods to generate multiple hyperplanes in the latent space using NN to predict pathloss will inevitably lead to poor performance. Furthermore, the self-supervised training method that uses data which is partially masked as input to train NN to predict masked information is mainly used in the training of generative models [24]. In the context of RM construction, the data representing environmental information, can be regarded as the RM whose pathloss is masked, and the NN-based RM construction methods use these masked RM data to recover the pathloss, which is a self-supervision training method. Based on the above analysis, RM construction is a generative problem, so from the perspective of the alignment of the problem and the methodology, a generative method should be used to effectively construct the RM. Moreover, the generative training methods enable NN training both using generative loss and MSE loss, where the NN’s capacity to extract intricate texture features of RM can be enhanced. Therefore, to achieve high-performance RM construction by taking advantage of the generative NN training method, in this paper, the RM construction problem is modeled as a conditional generative problem, where a generative diffusion-based method, named RadioDiff, is proposed to construct the RM effectively  [^2] In this paper, the terms “diffusion model” and “denoise diffusion model” are used interchangeably, both referring to methods for generating the required data from noise input. The diffusion process involves diffusing the data into noise, while the denoising process reconstructs the data from this noisy input.. The main contributions of this paper are as follows.
@@ II-B1 Forward Diffusion Procedure
where $\mathcal{N}(\cdot,\cdot)$ is Gaussian distribution, $\bm{I}$ is identity matrix, and $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$.
@@ IV-B RadioDiff via Decoupled Diffusion Model
where $\gamma_{t}$ and $\delta_{t}$ are hyper-parameters, and $\delta_{t}$ is designed to increase gradually over time while $\gamma_{t}$ decreases. According to [30], (8) can also be expressed as the following differential equation.
@@ IV-C Adaptive FFT Filter for DRM Enhancement
As shown in Fig. LABEL:fig-example, RM has lots of edge texture features, particularly in DRM, resulting in more high-frequency information in the frequency domain [32]. Although conventional convolutional layers within the denoising U-Net are effective in feature extraction, they face limitations in precisely extracting features from specific frequency components. Consequently, it is challenging for NN just based on conventional convolutional layers to achieve high performance, as they struggle to effectively extract high-frequency information caused by dynamic obstacles within the RM, since utilizing only traditional convolutional layer-based NN tends to construct an over-smooth RM, leading to blurred representations in the constructed RM. Therefore, as shown in Fig. LABEL:fig_method, an adaptive fast Fourier transform (FFT) filter (AFT) module is used to enhance the high-frequency features extracting capability of the diffusion model. AFT operates by initially transforming the 2D encoder-generated feature maps $\bm{z}$, with spatial dimensions $H\times W$ and channel count $C$, from the spatial domain to the frequency domain, represented as $\bm{z}_{c}=\mathcal{F}(\bm{z})$, where $\mathcal{F}$ denotes the FFT operation. Furthermore, AFT introduces a learnable weight matrix $\bm{w}\in\mathbb{C}^{H\times W\times C}$, which is applied to the frequency-domain features $\bm{z}_{c}$ using a Hadamard product, denoted as $\bm{w}\odot\bm{z}_{c}$. This learnable weight matrix enables the filter to dynamically adjust to the frequency distribution of the target data, thereby enhancing the model’s focus on relevant frequencies while attenuating less significant ones. This spectral filtering approach allows precise global frequency adjustments based on the importance derived from the data, ensuring that only relevant frequencies are amplified or attenuated. After the adaptive filtration process, the noise-reduced frequency domain signal is reconverted to the spatial domain using the Inverse Fast Fourier Transform (IFFT). To ensure the retention of valuable information and mitigate potential loss due to the filtering process, a residual connection from the raw feature map $\bm{z}$ to the output feature map is used. Thus, the operation of AFT can be formulated as follows.
@@ V-A Datasets
In this study, we evaluate the performance of the proposed method using the RadioMapSeer dataset provided by the pathloss RM construction challenge [33]. The dataset comprises 700 maps, each featuring 80 transmitter locations, accompanied by corresponding ground truths. The city maps cover cities such as Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv. In the dataset, both transmitter and receiver heights are set at 1.5 meters, while building heights are set at 25 meters. Each map is converted into a 256 × 256 pixels morphological 2D image with binary pixel values (0/1), where each pixel represents one square meter: ‘1’ inside buildings and ’0’ outside. Transmitter positions are stored in a two-dimensional numerical format and depicted in morphological images, with the transmitter’s pixel set to ‘1’ and all others to ‘0’. The transmitter power is set to 23 dBm, and the carrier frequency is 5.9 GHz. To obtain a sufficiently accurate RM as ground truth for training, the ground truth RM in the dataset is constructed using Maxwell’s equations, where the pathloss is calculated by the reflection and diffraction of electromagnetic rays. The RMs that only consider the impact of the static buildings on the electromagnetic rays are used as the ground truth of SRM. Additionally, the RMs in the dataset that consider the impact both of the static buildings and the vehicles, which are randomly generated along the roads, are used as the ground truth of the DRM.
@@ V-B Metrics
To comprehensively evaluate the quality of RM constructions, we utilize MSE, structural similarity index measurement (SSIM), and peak signal-to-noise ratio (PSNR) simultaneously as metrics, where MSE quantifies the overall error that is the accuracy of the RM construction, SSIM evaluates the preservation of structural information and detail precision, and PSNR measures the ratio of signal to noise affecting the fidelity of the constructed RM.
@@ V-B2 SSIM
SSIM is a quality assessment metric inspired by the human visual system. Since SSIM focuses on measuring texture differences, and there are lots of high-frequency details in RM, SSIM is suitable for evaluating the quality of the generated results. The SSIM metric evaluates three distinct components: luminance, contrast, and structural contrast, which can be calculated as follows.
@@ V-C Implementation
We implemented our RadioDiff framework using the PyTorch framework. The training process is divided into two phases, both utilizing an AdamW optimizer with a decaying learning rate, starting from $5\times 10^{-5}$ and reducing to $5\times 10^{-6}$. In the initial phase, the autoencoder is trained using RM images from the entire dataset’s training set as ground truth. This phase, which trains the VAE, takes approximately 120 hours on 4×4$\times$4 × NVIDIA A100 SXM GPUs with a batch size of 2. The subsequent phase involves training the denoise diffusion model, which takes around 360 hours using 4×4$\times$4 × NVIDIA A100 SXM GPUs with a batch size of 64. Although a larger batch size can achieve better training results and higher training speed [24], the batch size in the first stage is significantly smaller than that in the second stage. In the first stage, both the input and output are raw image data, which consumes substantial video memory. In contrast, during the second stage, when training the diffusion model, the data is processed in the latent space encoded by the VAE, thereby reducing video memory consumption. Consequently, a larger batch size can be employed to enhance training speed and accuracy. This highlights the importance of using the VAE to encode data into the latent space. For hyper-parameters, the $T$ is $500$, and $\bm{f}_{t}$ is $-\bm{z}_{0}$.
@@ V-D Comparisons with SOTA Methods
To evaluate the proposed RadioDiff model, we compare it with other SOTA methods. The following method is used to compare.
  • RadioUnet [17]: RadioUnet is one of the most effective sampling-free NN-based RM construction method, where a convolutional U-Net is used as the backbone NN, and the supervised learning is used to train the RadioUnet.
  • UVM-Net [34]: The training method and settings of UVM-Net are same as RadioUnet, but the backbone network is replaced by a convolutional layer with the latest attention-based Mamba layer.
  • RME-GAN [20]: The SOTA NN-based RM construction method, which uses a generative model cGAN to construct RM. However, RME-GAN is sampling-based, it not only uses environmental features but also uses the sampling pathloss as the input to construct the RM. For a fair comparison, the RME-GAN in this paper only uses environmental features as input.
@@ V-F Limitations and Discussion
Although this paper proposes a pioneer exploration of utilizing diffusion models for RM construction, and the proposed RadioDiff achieves SOTA performance, the issue of efficiency remains a key consideration for such a large generative model-based method. Table IV presents a comparison of the inference time and memory usage between the RadioDiff model and alternative models, showing that the diffusion model has higher resource demands during inference than other methods. However, it should be emphasized that, while the proposed RadioDiff is more time-consuming than other NN-based methods, the inferencing latency is still smaller than 1 second, which makes it still delay-acceptable for constructing RM dynamically. Moreover, techniques such as NN compression and efficient inference methods like denoising diffusion implicit models (DDIM) [35] can be leveraged to notably enhance efficiency, raising a trade-off between performance and efficiency, which stands as a promising direction for future research.
@@
TABLE IV: Inference Time and Memory Consuming.

### v2

@@ I Introduction
In wireless networks, pathloss quantifies the attenuation of signal strength between a pair of sender and receiver caused by free-space propagation loss and interactions of radio waves with obstacles [1, 2], which is critical to be measured for wireless resource management [3, 4, 5]. Traditionally, the pathloss measurement usually depends on pilot transmission and signal processing [6]. However, the dramatic increase of network nodes and antennas has led to the challenge of estimating high-dimensional channels [7], resulting in significant costs in training and feedback overhead, as well as signal processing complexity [8]. This issue further deteriorates in high-mobility scenarios with short channel coherence times for low-latency applications [9, 10]. Meanwhile, the upcoming 6G networks will introduce a large variety of node types, including passive equipment such as intelligent reflective surfaces (IRS) [11], which cannot actively transmit pilots or engage in digital signal processing to measure the pathloss.
@@
The fundamental reason for the above challenges is the misalignment between the generative problem attributes of the RM construction and the discriminant methodology. RM construction shows the characteristics of generative problems in both data features and NN training methods, while the existing sample-free NN-based methods use discriminant methods to construct RM, which inevitably limits the construction performance of RM. Specifically, in terms of data features, neither the value nor location of the elements of the pathloss to be predicted exist in the input environmental data, thus NN is required to generate pathloss from the raw data. Moreover, since the elements in the RM are not discrete values, it is almost impossible to construct the RM by dividing a finite number of hyperplanes to classify the elements in the environmental data as a discriminative problem [23]. Therefore, the use of discriminant training methods to generate multiple hyperplanes in the latent space using NN to predict pathloss will inevitably lead to poor performance. Furthermore, the self-supervised training method that uses data which is partially masked as input to train NN to predict masked information is mainly used in the training of generative models [24]. In the context of RM construction, the data representing environmental information, can be regarded as the RM whose pathloss is masked, and the NN-based RM construction methods use these masked RM data to recover the pathloss, which is a self-supervision training method. Based on the above analysis, RM construction is a generative problem, so from the perspective of the alignment of the problem and the methodology, a generative method should be used to effectively construct the RM. Moreover, the generative training methods enable NN training both using generative loss and MSE loss, where the NN’s capacity to extract intricate texture features of RM can be enhanced. Generative adversarial networks (GANs) have been extensively studied in RM reconstruction; however, their potential is often hindered by extreme instability during training. In contrast, the diffusion model offers a novel solution to the RM reconstruction problem. Geographic maps and RM contain abundant high-frequency sharp edges. Directly feeding these into NN-based models can result in erratic predictions, noise, and artifacts. Diffusion models, however, demonstrate superior capability in capturing this edge information and predicting conditions. Therefore, to achieve high-performance RM construction by taking advantage of the generative NN training method, in this paper, the RM construction problem is modeled as a conditional generative problem, where a generative diffusion-based method, named RadioDiff, is proposed to construct the RM effectively  [^2] In this paper, the terms “diffusion model” and “denoise diffusion model” are used interchangeably, both referring to methods for generating the required data from noise input. The diffusion process involves diffusing the data into noise, while the denoising process reconstructs the data from this noisy input.. The main contributions of this paper are as follows.
@@ II-B1 Forward Diffusion Procedure
where $\mathcal{N}(\cdot,\cdot)$ is Gaussian distribution, $\bm{I}$ is identity matrix, and $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$. In addition, Eq. (4) describes how the noisy data $\bm{x}_{t}$ is generated by combining the original input $\bm{x}_{0}$ with a Gaussian noise term $\bm{\epsilon}$. The term $\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}$ represents the scaled contribution of the original data, while $\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ represents the amount of noise added at each step. As $t$ increases, the influence of the noise term grows, resulting in progressively noisier data.
@@ IV-B RadioDiff via Decoupled Diffusion Model
where $\gamma_{t}$ and $\delta_{t}$ are hyper-parameters, and $\delta_{t}$ is designed to increase gradually over time while $\gamma_{t}$ decreases. Compared to traditional diffusion models where the diffusion process directly adds noise to the original input $\bm{x}_{0}$ or $\bm{z}_{0}$, the decoupled diffusion model used in RadioDiff splits this process into two distinct stages. In DDM, $\bm{z}_{0}$ first diffuses into a $\bm{0}$ vector, which effectively decouples the contribution of the original input from the added noise. This allows for a more controlled diffusion process, as the noise $\bm{\epsilon}$ is only introduced after $\bm{z}_{0}$ has been reduced to a $\bm{0}$ vector. As a result, DDM reduces the variance in early diffusion steps and improves stability during both training and inference. Furthermore, this decoupled structure helps mitigate the prolonged inference time seen in traditional diffusion models, enabling more efficient generation in RadioDiff. Thus, according to [30], (8) can also be expressed as the following differential equation.
@@ IV-C Adaptive FFT Filter for DRM Enhancement
As shown in Fig. LABEL:fig-example, RM exhibits numerous edge texture features, especially in DRM, which generate substantial high-frequency information in the frequency domain [32]. Although conventional convolutional layers within the denoising U-Net effectively extract features, they struggle to precisely capture high-frequency components . As a result, neural networks (NNs) based solely on traditional convolutional layers often produce overly smooth RM outputs, leading to blurred representations and suboptimal performance in dynamic environments.
To address this issue, as depicted in Fig. LABEL:fig_method, an adaptive Fast Fourier Transform (FFT) filter (AFT) module is introduced, specifically to enhance the model’s capacity for extracting high-frequency features. The AFT module operates by transforming the 2D feature maps $\bm{z}$, generated by the encoder with spatial dimensions $H\times W$ and channel count $C$, from the spatial domain to the frequency domain via the FFT operation, represented as $\bm{z}_{c}=\mathcal{F}(\bm{z})$, where $\mathcal{F}$ denotes the FFT. To improve the model’s ability to focus on relevant frequency components, AFT incorporates a learnable weight matrix $\bm{w}\in\mathbb{C}^{H\times W\times C}$. This matrix is applied to the frequency-domain features $\bm{z}_{c}$ via the Hadamard product, $\bm{w}\odot\bm{z}_{c}$. The learnable weight matrix dynamically adjusts the model’s response to different frequency distributions in the target data, emphasizing important frequencies while attenuating irrelevant ones . This adaptive spectral filtering allows the model to perform global frequency adjustments based on data-driven relevance, ensuring that high-frequency features are enhanced while less critical frequencies are suppressed .
Following this adaptive filtering process, the modified frequency-domain features are transformed back into the spatial domain using the Inverse Fast Fourier Transform (IFFT). To preserve key information and mitigate potential losses during filtering, a residual connection is introduced between the original feature map $\bm{z}$ and the output feature map . The complete operation of the AFT module can be described as follows:
@@ V-A Datasets
In this study, we evaluate the performance of the proposed method using the RadioMapSeer dataset provided by the pathloss RM construction challenge [33]. The dataset consists of 700 maps, each with unique geographic information (e.g., building data), with each map containing 80 transmitter locations and their corresponding ground truth data. Each map contains between 50 and 150 buildings. We selected 500 maps for the training dataset and the remaining 200 maps for the test dataset. There is no overlapping terrain information between the training and test datasets.
The city maps include data from cities such as Ankara, Berlin, Glasgow, Ljubljana, London, and Tel Aviv, sourced from OpenStreetMap. In the dataset, both transmitter and receiver heights are set at 1.5 meters, while building heights are set at 25 meters. Each map is converted into a 256 × 256 pixel morphological 2D image with binary pixel values (0/1), where each pixel represents one square meter: ‘1’ for areas inside buildings and ‘0’ for areas outside. Transmitter positions are stored in a two-dimensional numerical format and depicted in morphological images, with the transmitter’s pixel set to ‘1’ and all others to ‘0’. The transmitter power is set to 23 dBm, and the carrier frequency is 5.9 GHz. To obtain a sufficiently accurate RM as ground truth for training, the ground truth RM in the dataset is constructed using Maxwell’s equations, where the pathloss is calculated by the reflection and diffraction of electromagnetic rays. The RMs that only consider the impact of the static buildings on the electromagnetic rays are used as the ground truth of SRM. Additionally, the RMs in the dataset that consider the impact both of the static buildings and the vehicles, which are randomly generated along the roads, are used as the ground truth of the DRM, as shown in Fig. LABEL:fig-example.
@@ V-B Metrics
To comprehensively evaluate the quality of RM construction, we begin by adopting the parameters commonly used in previous studies [17], namely NMSE and RMSE. Additionally, we observe that the accurate generation of structural information and details is a key objective in RM reconstruction tasks, whereas the MSE index focuses on overall error and does not directly address these specific requirements. Therefore, we propose introducing structural similarity index measurement (SSIM) and peak signal-to-noise ratio (PSNR) as additional metrics in this paper. SSIM evaluates the preservation of structural information to emphasize the accuracy of structural detail reconstruction, while PSNR measures the signal-to-noise ratio to assess the fidelity of RM construction, particularly with respect to edge signal reconstruction.
@@ V-B2 SSIM
SSIM is a quality assessment metric inspired by the human visual system. Since SSIM focuses on measuring texture differences, and there are lots of high-frequency details in RM, SSIM is suitable for evaluating the quality of the generated results. We also believe that greater attention should be given to the brightness of the signal radiation, the contrast between the signal radiation and the surrounding area, and the accuracy of geographic map in RM reconstruction. This aligns with the SSIM metric, which evaluates three key components: brightness, contrast, and structural information , which can be calculated as follows.
@@ V-C Implementation
We implemented our RadioDiff framework using the PyTorch framework. The training process is divided into two phases, both utilizing an AdamW optimizer with a decaying learning rate, starting from $5\times 10^{-5}$ and reducing to $5\times 10^{-6}$. In the initial phase, the autoencoder is trained using RM images from the entire dataset’s training set as ground truth. This phase, which trains the VAE with $z$-channels set to 3 and embedding dimension of 128, takes approximately 120 hours on 4×4$\times$4 × NVIDIA A100 SXM GPUs with a batch size of 2. The subsequent phase involves training the denoise diffusion U-net model, which takes around 360 hours using 4×4$\times$4 × NVIDIA A100 SXM GPUs with a batch size of 64. In this implementation, the input image size is $256\times 256$. The diffusion process uses $T=1000$ in training, and the loss function is $l2$-based with an objective of predicting KC. The start distribution is set to normal, and the perceptual weight is set to 00. Although a larger batch size can achieve better training results and higher training speed [24], the batch size in the first stage is significantly smaller than that in the second stage. In the first stage, both the input and output are raw image data, which consumes substantial video memory. In contrast, during the second stage, when training the diffusion model, the data is processed in the latent space encoded by the VAE, thereby reducing video memory consumption. Consequently, a larger batch size can be employed to enhance training speed and accuracy. This highlights the importance of using the VAE to encode data into the latent space. The hyper-parameter for the diffusion timesteps $T$ is set to 500 for inference, and the diffusion process is decoupled with $\bm{f}_{t}$ defined as $-\bm{z}_{0}$.
@@ V-D Comparisons with SOTA Methods
To evaluate the proposed RadioDiff model, we compare it with other SOTA methods. To ensure a comprehensive comparison of the experiments, we compare the CNN-based, GAN-based, and Mamba-based methods separately, which represent the primary architectures used in the current RM reconstruction task based on deep learning. For the detailed parameter settings of the comparison model, we adhere to the description provided in the article [17, 34, 20]. Additionally, for fairness, the training and test data will be aligned with RadioDiff. The following method is used to compare.
  • RadioUnet [17]: RadioUnet is one of the most effective sampling-free NN-based RM construction method, where a convolutional U-Net is used as the backbone NN, and the supervised learning is used to train the RadioUnet. RadioUnet reconstructs the wireless propagation graph using a simple yet effective network architecture that learns environmental characteristics, making it one of the most representative reconfiguration algorithms based on reel machines.
  • UVM-Net [34]: The training method and settings of UVM-Net are same as RadioUnet, but the backbone network is replaced by a convolutional layer with the latest SSMs (State Space Sequence Models). SSMs are specifically designed to handle long-sequence data and are particularly well-suited for modeling long-range dependencies. SSMs map the input sequence to a hidden state through a state space model and predict the output based on this hidden state. As a result, SSMs exhibit enhanced local feature capture and efficient remote modeling capabilities. We selected UVM-Net as the baseline to evaluate the performance differences between this SSM-driven architecture and traditional convolutional networks in the wireless propagation graph reconstruction task.
  • RME-GAN [20]: The SOTA NN-based RM construction method, which uses a generative model cGAN to construct RM. However, RME-GAN is sampling-based, it not only uses environmental features but also uses the sampling pathloss as the input to construct the RM. For a fair comparison, the RME-GAN in this paper only uses environmental features as input. As GAN models have been at the forefront of generative model research in recent years, they effectively showcase the unique capabilities and challenges of generative adversarial networks in handling RM reconstruction tasks.
@@ V-F Limitations and Discussion
Although this paper proposes a pioneer exploration of utilizing diffusion models for RM construction, and the proposed RadioDiff achieves SOTA performance, the issue of efficiency remains a key consideration for such a large generative model-based method. Table IV presents a comparison of the inference time and memory usage between the RadioDiff model and alternative models, showing that the diffusion model requires more resources and has a longer inference time than other methods. However, it should be emphasized that although the proposed RadioDiff is more time-consuming than other NN-based methods, the inference delay is still less than one second, which remains acceptable for dynamically constructing RM. For training data, RadioDiff uses the same dataset as other NN-based methods, and no limitations regarding its access to training data were identified in the experiment. In addition, although VAE needs to be trained separately before RadioDiff’s formal training, this is due to LDM being used for the RM reconstruction task for the first time. In subsequent research, the relevant pre-trained weights of VAE can be directly used without re-training, which will significantly streamline future research efforts. Moreover, techniques such as NN compression and efficient inference methods like denoising diffusion implicit models (DDIM) [35] can be leveraged to notably enhance efficiency, raising a trade-off between performance and efficiency, which stands as a promising direction for future research.
@@
TABLE IV: Inference Time and Memory Consuming. Results in bold red and underlined blue highlight the highest and second highest, respectively.